OpenStudy (anonymous):

what is the integral of 1/(16x^2-1)

OpenStudy (anonymous):

Let \(x=\dfrac{1}{4}\sec u\), so that \(dx=\dfrac{1}{4}\sec u\tan u~du\). \[\int\frac{1}{16x^2-1}~dx=\int\frac{\dfrac{1}{4}\sec u\tan u}{16\left(\dfrac{1}{4}\sec u\right)^2-1}~du=\frac{1}{4}\int\frac{\sec u\tan u}{\sec^2 u-1}~du\]

OpenStudy (dumbcow):

another approach is to split it into 2 fractions \[\frac{1}{16x^{2} -1} = \frac{1}{(4x+1)(4x-1)} = \frac{A}{4x+1} + \frac{B}{4x-1}\] where \[A(4x-1) + B(4x+1) = 1\] 4A +4B = 0 --> A = -B B -A = 1 --> B = 1/2 --> A = -1/2 integral becomes \[\frac{1}{2} \int\limits \frac{1}{4x-1} - \frac{1}{4x+1} dx\] use \[\int\limits \frac{1}{u} = \ln u\]